10th CONGRESS OF YUGOSLAV MATHEMATICIANS

Belgrade, January 21–24, 2001

The 10th Congress of Yugoslav Mathematicians was held January 21 to 24, 2001, in Belgrade. The organizers were: Faculty of Mathematics in Belgrade, Mathematical Society of Serbia, Union of Mathematical Societies of Yugoslavia and Institute of Mathematics of the Faculty of Sciences in Novi Sad.

Previous congresses were: I congress (1949) – Bled; II congress (1955) – Zagreb; III congress (1960) – Beograd; IV congress (1965) – Sarajevo; V congress (1970) – Ohrid; VI congress (1975) – Novi Sad; VII congress (1980) – Bečići; VIII congress (1985) – Priština; IX congress (1995) – Petrovac.

The opening ceremony was completed with two plenary lectures:

Stevo Todorčević: Hierarchies of Ackermann and Baire;

Judita Cofman: How to motivate 10–18-year old students to solve various mathematical problems on their own.

During the later work of the Congress, 14 invited lectures (Siniša Crvenković, Dragoš Cvetković, Milutin Dostanić, Vladimir Dragović, Kosta Došen, Aleksandar Ivić, Danko Jocić, Nenad Mladenović, Nedeljko Parezanović, -Dura Paunić, Slaviša Prešić, Vladimir Rakočević, Rade Živaljević, Miodrag Živković) as well as 181 section lectures in 12 Sections were delivered. Here are the titles of the Sections, together with the number of lectures in each of them: Logic (7), Algebra (15), Geometry and Topology (15), Real and Complex Analysis (16), Functional Analysis and Operator Theory (25), Differential Equations (8), Combinatorics and Graph Theory (2), Probability and Statistics (11), Numerical Analysis and Optimization (10), Computer Science (17), Applied Mathematics (26), Teaching, History and Popularizing of Mathematics (29).

The Republic Seminar 2001 on Mathematics and Computer Science Education in primary and secondary schools, and universities was the part of the Teaching section of the Congress.

There were about nine hundred participating mathematicians, including thirty guests from Austria, Italy, FYR Macedonia, Germany, Nigeria, Romania and USA. Four round tables were organized; their titles reflect the discussed topics:

• Mathematics Instruction and reform of the educational system;

• Where is Theoretical Mathematics going?;

• Industrial Mathematics and Informatics;

• Mathematics and mathematicians today and tomorrow.

A prize was awarded by the Organizing Committee to Danko Jocić form the University of Belgrade for the contributions in mathematical sciences by researchers up to 40 years old.

The regular Assembly of the Union of Mathematical Societies of Yugoslavia was held during the Congress. The new Executive Board of the Union was elected, Zoran Kadelburg from the University of Belgrade being its new president. Also, Mathematical Society of Serbia held its Assembly and Rade Doroslovački from the University of Novi Sad was elected as the new president of the Society.

The Proceedings of the 10th Congress of Yugoslav Mathematicians were published by the Faculty of Mathematics in Belgrade.

Organizing Committee